OpenStudy (anonymous):

Help with two problems please! Write an equation in standard form for a line with an x intercept of 2 and a y intercept of 5. Write the standard form of the equation of the line parallel to the graph of 2y-6 = 0 and passing through B (4, -1). Can someone help me solve these please? :)

OpenStudy (mathmate):

Use the symmetric form when both intercepts are given: \(\dfrac{x}{A}+\dfrac{y}{B}=1\) where A and B are the x and y intercepts respectively. Example: Line with x-intercept of 4 and y-intercept of 5: \(\dfrac{x}{4}+\dfrac{y}{5}=1\) Multiply by 20 on both sides: 5x+4y=20, or in standard form: 5x+4y-20=0

OpenStudy (anonymous):

this is for the question im assuming, right?

OpenStudy (anonymous):

i got 5x + 2y - 10 = 0...does that sound ok?

OpenStudy (anonymous):

for the first question i meant! haha

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

yes, 5x + 2y - 10 = 0 looks good for the first one.

OpenStudy (anonymous):

thanks so much! :) @mathmate

OpenStudy (mathmate):

You're welcome!

OpenStudy (mathmate):

Do you need help with the second one?